# When Can Transformers Count to n?

### v1

@@ 5.3.1 Limitations of the CountAttend solution
First, Proposition 5.3 has an $O(n)$ sized MLP, which is a result of its internal implementation of the function $x\mapsto\frac{1}{x}$ using an MLP, where $x\in\left[\frac{1}{n},1\right]$, and the desired precision is $0.5$ (because we are counting, and can round the result). In the proof of 5.3 we used a naive implementation of this function. It is natural to ask if a smaller implementation exists. The following result shows this is not possible.
@@ 6.1 MFE Models Must have $d\geq\Omega(m)$
We begin by showing a lower bound on the required size of a $1$-layer transformer solving the MFE task. The following results establishes that MFE can be implemented only when $dhp=\Omega(\min\{m,n\})$. This means that at a fixed precision $p$ and if $n>m$, the dimension $d$ must grow linearly in the vocabolary size in order to implement MFE. This is the same trend that we saw for the QC problem.
@@
The full proof can be found in Appendix B. The proof uses a communication complexity argument that is inspired by a lower bound of Sanford et al. [12]. This lower bound can be interpreted as if transformers that solve the MFE task need to have an embedding size (i.e. dimension or number of precision bits) that scales with the size of the dictionary, or have many attention heads. Note that increasing the size of the MLP which follows the attention cannot break the lower bound.
@@ 6.2 MFE Can be Implemented when $d=O(m)$
One limitation of the above result is that it requires an MLP that grows with $m$. This can be avoided if using two layers of attention. A two layer implementation is simple: use the first layer to calculate the “Query Count” for each element, and then use softmax to calculate the maximum over tokens. This construction does not need an MLP at all. Another option is to use a depth $\log(m)$ MLP with $m$ neurons at each layer to calculate the maximum from the histogram (see [11]), however having depth which relies even logarithmically on the dictionary size is infeasible for practical implementations.

### v2

@@ 4.4.1 Limitations of the CountAttend solution
First, Proposition 4.3 has an $O(n)$ sized MLP, which is a result of its internal implementation of the function $x\mapsto\frac{1}{x}$ using an MLP, where $x\in\left[\frac{1}{n},1\right]$, and the desired precision is $0.5$ (because we are counting, and can round the result). In the proof of 4.3 we used a naive implementation of this function. It is natural to ask if a smaller implementation exists. The following result shows this is not possible.
@@ 5.1 MFE Models Must have $d\geq\Omega(m)$
We begin by showing a lower bound on the required size of a $1$-layer transformer solving the MFE task. The following result establishes that MFE can be implemented only when $dhp=\Omega(\min\{m,n\})$. This means that at a fixed precision $p$ and if $n>m$, the dimension $d$ must grow linearly in the vocabolary size in order to implement MFE. This is the same trend that we saw for the QC problem.
@@
The full proof can be found in Appendix B. The proof uses a communication complexity argument that is inspired by a lower bound of Sanford et al. [13]. This lower bound implies that transformers which solve the MFE task need to have an embedding size that scales with the size of the dictionary, or have many attention heads. Note that increasing the size of the MLP which follows the attention cannot break the lower bound.
@@ 5.2 MFE Can be Implemented when $d=O(m)$
One limitation of the above result is that it requires an MLP that grows with $m$. This can be avoided if using two layers of attention. A two layer implementation is simple: use the first layer to calculate the “Query Count” for each element, and then use softmax to calculate the maximum over tokens. This construction does not need an MLP at all. Another option is to use a depth $\log(m)$ MLP with $m$ neurons at each layer to calculate the maximum from the histogram [see 12], however having depth which relies even logarithmically on the dictionary size is infeasible for practical implementations.
@@ Appendix C Inner Product of Random Vectors
Figure 3: Evaluation of Gemini 1.5 on the MFC task. See Sec. D.
Appendix D Empirical Results on the MFC Task
In Sec. LABEL:sec:gemini we provided results for running Gemini 1.5 on the QC task. Here we provide results for the MFC task. Experimental setting is similar to QC, with the samples drawn in the same way. The prompt has the following format: “You are given an array of length 5. Find the count of the most frequent element in the array. The array is [1,1,2,2,3]. How many times does the most frequent word in the array appear? Respond in just one number. No additional text.”. Another difference from QC was that the sequence length was set such that the expected MFC would be 40, for the given vocabulary size $m$. This was done via simulation (since no closed form expression is available for expected MFC). For the binary baseline in this case we cannot choose an MFC task, because in the binary case the MFC will be larger than our desired expected count (i.e., 40). Instead, we take the binary baseline to be the minimum frequency count, with expected minimum equal to $40$. Results are provided in Fig. 3, and show that results deteriorate with vocabulary size (also with respect to the binary baseline) again suggesting that vocabulary size affects the complexity of this task, as our theoretical results suggest.